OpenStudy (anonymous):

There are initially 1000 bacteria in a culture. The number of bacteria doubles each hour. The number of bacteria N present after t hours is . How long will it take the culture to increase to 50,000 bacteria?

OpenStudy (anonymous):

solve for \(t\), \[50,000=1000\times 2^t\]

OpenStudy (anonymous):

the formula they gave me to use was n=1000(2)t

OpenStudy (anonymous):

two steps only divide by 1,000 get \[50=2^t\] then change of base gives \[t=\frac{\ln(50)}{\ln(2)}\]

OpenStudy (anonymous):

5.6 got it thank you :)